5. Suppose X and Y are standard normal random variables. Find an expres- sion for P(X - 3Y S1) in terms of the standard normal distribution function In two cases: (i) X and Y are independent (ii) X and Y have bivariate normal distribution with correlation ρ-1/2.

please help me
Answer #1

TOPIC:Application of the Normal distribution.

5.) (i) Griven that, XNN (0,D. YN N(0,0. X independently, Then, (X-3D) is a linean combination of the independent Nommal variCry Here, X and Y are bivariate Nommal, with, S=Conn. (x, y) = % . i. Si Cov (x,x) a) sd (X). sd (y) = { sd(x). So cor (x: 1)= var (x) + 9 var (V) - 2×3 Cov (8,8). = van (x) + 9 var (y) – 6 Cov (8,9). = 1+ 9x1 - 672 = 7: i. (x-34) NN (0, 7). .. we ne